OpenStudy (anonymous):

I need to get this done! Hurry! What is the slope of the line that passes through the given points? (6, 2) and (7, 4)

OpenStudy (anonymous):

\[m = \frac{y_1 - y_2}{x_1 - x_2} \implies m = \frac{4 - 2}{7 - 6} \implies m = ?\]

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

Yup :)